5) Find the Fourier transform of exp(-ax) using the exponential form of the transform equations.
Answer #1

PLEASE RATE IT UP!!!!!! THANKS!!!!!!

f(1) = e-ar? F(R) = 8° f(0) e ikxdx= artik ndx -7 °6-a(z+- 2 x + (iki) - (1994) ___ ATT Jan s e-4 (-tak = p e-a (nt - - hot